OpenStudy (anonymous):

Is my answer correct? Question: Find the exact value of the following trig function: cos (-150deegrees) answer: Theta terminates in Quadrant III. Reference angle is 30degrees cos = negative in Quadrant III cos (-150) = -cos (30) = Square root of 3/2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Correct

OpenStudy (angela210793):

tht's right

OpenStudy (anonymous):

I have a q, If at the end its -cos (30) shouldnt my answer be influenced and become - Square root of 3/2? Or it doesnt matter?

OpenStudy (anonymous):

\[\cos (-\theta) = \cos \theta \]

OpenStudy (anonymous):

Please dont leave, I have one more question:

OpenStudy (anonymous):

to find exact value of tan (225) is my answer correct? theta terminates in Q3 R. angle is 45 tan = positive in Q3 tan(225) = tan(45) = 1

OpenStudy (experimentx):

tan(180+45)

OpenStudy (experimentx):

whenever there is 180 turn function remains same, and or 90 function changes.

OpenStudy (angela210793):

i wish u knew albanian -_- i'd have explained these to u -_-

OpenStudy (anonymous):

Wait, so 1 is wrong?

OpenStudy (anonymous):

so exact balue of tan(225) is not 1?

OpenStudy (anonymous):

*value

OpenStudy (anonymous):

1 is right!

OpenStudy (anonymous):

Oh. lol. Ok thank you guys so much! as usual you guys ROCK!!:)